OpenStudy (anonymous):

How to find points of intersection between y = x^2 and y = -x^2 + 8?

OpenStudy (anonymous):

\[y = x^2\] \[y = -x^2 + 8\]

OpenStudy (lgbasallote):

equate them to one another \[x^2 = -x^2 + 8\] then solve for x do you need more help?

OpenStudy (anonymous):

Thank you! It also says to create a table of values for each equation. How is that done?

OpenStudy (lgbasallote):

im thinking you just do something like this  in the drawing the equation is y = x^2 then, i assigned values for x and solved for y when x = 1, y = 1 when x = 2, y = 2^2 = 4 when x = 3, y = 3^2 = 9 and so on... got it?